OpenStudy (anonymous):

A computer is infected with the Sasser virus. Assume that it infects 20 other computers within 5 minutes; and that these PCs and servers each infect 20 more machines within another five minutes, etc. How long until 100 million computers are infected? i dont get this stuff

OpenStudy (anonymous):

Lets try this, we are going to construct a table of pairsand see what comes out

OpenStudy (anonymous):

\[100,000,000=20^{\frac{t}{5}}\]

OpenStudy (anonymous):

\[\log(100,000,000)=\frac{t}{5}\log(20)\] \[8=\frac{t}{5}\log(20)\] \[t=\frac{40}{\log(20)}\]

OpenStudy (anonymous):

wat part is the answer?

OpenStudy (anonymous):

i got this 100000000 = 1*(20)^(t/5). Log both sides. log (100000000) = (t/5)*log(20). t = 5*log(100000000)/log(20) t = 30.75 minutes.